# Failure of uniform laws of large numbers for subdifferentials and beyond

### v1

@@ 3.3 Proof of Proposition 1
Since $f(\xi,\cdot)$ is real-valued and convex, by [13, Theorem 23.1], the directional derivative of $f(\xi,\cdot)$ exists at any point $x$ and in every direction $w$, and we denote it by $f_{x}^{\prime}(\xi,x;w)$. Using [20, Theorem 7.46], the function $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is convex and directionally differentiable with $\mathcal{A}$-measurable $f_{x}^{\prime}(\cdot,x;w)$ and $(\mathbb{E}[f(\bm{\xi},\cdot)])^{\prime}(x;w)=\mathbb{E}[f_{x}^{\prime}(\bm{\xi},x;w)].$ From [6, Theorems V.3.3.8, V.3.3.3], [13, Theorem 23.2], and the non-emptiness of compact $\partial_{x}f(\xi,x)$, we can write the Hausdorff distance as follows
${d\kern-1.49994ptl}\Big(\partial\mathbb{E}[f(\bm{\xi},\cdot)](x),\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x)\Big)=\sup_{|w|=1}\Big|\mathbb{E}[f_{x}^{\prime}(\bm{\xi},x;w)]-\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f_{x}^{\prime}(\bm{\xi}^{i},x;w)\Big|.$
Let $P=\mathbb{P}\circ(\bm{\xi}^{1})^{-1}$. Hence, it suffices to show that, for any fixed $w\in\{-1,1\}$, the class of $\mathcal{A}$-measurable functions $\mathcal{S}=\{f_{x}^{\prime}(\cdot,x;w)\mid x\in\mathbb{R}\}$ is $P$-Glivenko–Cantelli; i.e.,
@@
which is well defined as explained below. For simplicity, assume $w=1$. Using the fact that $f(\xi,\cdot)$ is $L(\xi)$-Lipschitz, we have $|f_{x}^{\prime}(\xi,x;w)|\leq L(\xi)$ for any $x\in\mathbb{R}$. Because $\mathbb{E}[L(\bm{\xi})]<\infty$, the class $\mathcal{S}$ admits an integrable envelope $L$. From [20, Theorem 7.42] and [13, Theorem 24.1], the function $t\mapsto f_{x}^{\prime}(\xi,t;w)$ is non-decreasing and right-continuous (when w=1w=-1, it is non-increasing and left-continuous), thereby $\mathcal{S}$ is pointwise measurable [22, Example 2.3.4] and hence $P$-measurable [22, Definition 2.3.3] with (7) being well defined. With monotonicity and boundedness of $f_{x}^{\prime}(\xi,\cdot;w)$, set
@@
where $l,u:\Xi\to\mathbb{R}$ are $\mathcal{A}$-measurable with $\mathbb{E}[u(\bm{\xi})]-\mathbb{E}[l(\bm{\xi})]\leq 2\mathbb{E}[L(\bm{\xi})]$. If $\mathbb{E}[u(\bm{\xi})]=\mathbb{E}[l(\bm{\xi})]$, then $f_{x}^{\prime}(\xi,\cdot;w)$ is almost surely constant and the claim is trivial. Assume $\mathbb{E}[(u-l)(\bm{\xi})]>0$, and fix any $0<\varepsilon<\tfrac{2}{3}\mathbb{E}[(u-l)(\bm{\xi})]$. For each $n\in\mathbb{N}$, define
@@
where the attainment is from the right-continuity of $t\mapsto\mathbb{E}[f_{x}^{\prime}(\bm{\xi},t;w)]$. Let $N\in\mathbb{N}$ be the first number such that $\mathbb{E}[f_{x}^{\prime}(\bm{\xi},t_{N};w)]\geq\mathbb{E}[u(\bm{\xi})]-\tfrac{\varepsilon}{2}$. By construction, we have $N\leq 1+2\mathbb{E}[L(\bm{\xi})]/\varepsilon<\infty$ and $\mathbb{E}[|u(\bm{\xi})-f_{x}^{\prime}(\bm{\xi},t_{N};w)|]\leq\varepsilon$. By [13, Theorem 24.1], we get $f_{x}^{\prime}(\xi,t_{n};-w)=\lim_{s\uparrow t_{n}}f_{x}^{\prime}(\xi,s;w)$. Using monotone convergence theorem and the definition of $t_{n}$, for any $1\leq n<N$, one has
@@
which yields that $\mathbb{E}[|-f_{x}^{\prime}(\bm{\xi},t_{n+1};-w)-f_{x}^{\prime}(\bm{\xi},t_{n};w)|]\leq\varepsilon$ and $\mathbb{E}[|-f_{x}^{\prime}(\bm{\xi},t_{1};-w)-l(\bm{\xi})|]\leq\varepsilon$. Then, the following sets of functions
@@
Hence, the bracketing number $\mathcal{N}_{[\,]}(\varepsilon,\mathcal{S},L^{1}(P))$ is finite. Invoking [22, Theorem 2.4.1], we conclude that the class $\mathcal{S}$ is $P$-Glivenko–Cantelli. Applying the same argument to w=1w=-1, with simple adjustments for nonincreasingness and left-continuity, and then taking the maximum over $w\in\{-1,1\}$ completes the proof.
@@ 3.4 Proof of Theorem 3
$f(\xi,x)=h(g(\xi,x))+35\|x\|^{2}=\max\{g(\xi,x),0\}+35\|x\|^{2}.$
Define $\delta^{\nu}=(2240\lceil 2^{\nu+1}\log(\nu+1)\rceil^{4})^{-1}.$
@@
$\|g(\xi,x)-g(\xi,y)\|\leq|x_{1}-y_{1}|+\mathop{\sum}\nolimits_{j\in\{k_{1},k_{2}\}}|\psi_{j}(x_{2})-\psi_{j}(y_{2})|\leq 61\|x-y\|.$
Then, $g(\xi,\cdot)$ is $C$-Lipschitz continuous for any $\xi\in\Xi$ with
$\nabla_{x}g(\xi,x)=\left[\begin{array}[]{c}1\\ \mathop{\sum}\nolimits_{k=1}^{\infty}\psi_{k}^{\prime}(x_{2})(2\mathop{\textnormal{bit}}\nolimits_{k}(\xi)-1)\end{array}\right].$
Similarly, one has $\|\nabla_{x}g(\xi,x)-\nabla_{x}g(\xi,y)\|\leq C\|x-y\|$, so that the function $g(\xi,\cdot)$ is $C$-smooth. Hence, $f(\xi,\cdot)$ is convex and $2C$-Lipschitz on $X$. From Lemma 1, the functions $\xi\mapsto\mathop{\textnormal{bit}}\nolimits_{k}(\xi)$, $\xi\mapsto g(\xi,x)$, and then $\xi\mapsto f(\xi,x)$ are measurable. Therefore, $f$ and $g$ are Carathéodory. Moreover, for any $x\in\mathbb{R}^{2}$, one has $|\mathbb{E}[f(\bm{\xi},x)]|\leq\mathbb{E}[|g(\bm{\xi},x)|]+35\|x\|^{2}\leq|x_{1}|+\tfrac{1}{64}+35\|x\|^{2}<\infty.$
@@
and $f(\xi,\cdot)$ is smooth on $\cup_{k}\mathbb{B}(p_{k},\Delta_{k})$ for any $\xi\in\Xi$. Hence, for any $k\in\mathbb{N}$ and $y\in\mathbb{B}(p_{k},\Delta_{k})$, by [14, Exercise 10.26], we can write
@@
Let $\mathcal{F}$-measurable $E\subset\Omega$ with $\mathbb{P}(E)=1$, $\bar{\bm{\nu}}$, and $\{\bm{k}^{\nu}\}_{\nu}$ be defined in Lemma 2. For any $\omega\in E$, $\nu\geq\bar{\bm{\nu}}(\omega)$, one has $\omega\in E^{\nu}_{\bm{k}^{\nu}(\omega)}$ for some $\bm{k}^{\nu}(\omega)\leq\lceil 2^{\nu+1}\log(\nu+1)\rceil$. Then, $\mathop{\textnormal{bit}}\nolimits_{\bm{k}^{\nu}(\omega)}(\bm{\xi}^{i}(\omega))=1$ for all $i\in[\nu]$. Let $\bm{p}^{\nu}(\omega)=p_{\bm{k}^{\nu}(\omega)}\in\{p_{k}\}_{k}$. For any $y\in\mathbb{B}(\bm{p}^{\nu}(\omega),\Delta_{\bm{k}^{\nu}(\omega)})$, we have
@@
$0\leq\delta^{\nu}=\frac{1}{2240\lceil 2^{\nu+1}\log(\nu+1)\rceil^{4}}=\Delta_{\lceil 2^{\nu+1}\log(\nu+1)\rceil}\leq\Delta_{\bm{k}^{\nu}(\omega)}.$
Therefore, for any $\omega\in E$, we conclude that $\bm{p}^{\nu}(\omega)\in\{p_{k}\mid 1\leq k\leq\lceil 2^{\nu+1}\log(\nu+1)\rceil\}\subset D^{\nu}$ and

### v2

@@ 3.3 Proof of Proposition 1
Since $f(\xi,\cdot)$ is real-valued and convex, by [18, Theorem 23.1], the directional derivative of $f(\xi,\cdot)$ exists at any point $x$ and in every direction $w$, and we denote it by $f_{x}^{\prime}(\xi,x;w)$. Using [25, Theorem 7.46], the function $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is convex and directionally differentiable with Borel measurable $f_{x}^{\prime}(\cdot,x;w)$ and $(\mathbb{E}[f(\bm{\xi},\cdot)])^{\prime}(x;w)=\mathbb{E}[f_{x}^{\prime}(\bm{\xi},x;w)].$ By [7, Theorem 2.7.2], one has $\mathbb{E}[\partial_{x}f(\bm{\xi},x)]=\partial\mathbb{E}[f(\bm{\xi},\cdot)](x)$. From [10, Theorems V.3.3.8, V.3.3.3], [18, Theorem 23.2], and the non-emptiness of compact $\partial_{x}f(\xi,x)$, we can write the Hausdorff distance as follows
${d\kern-1.49994ptl}\Big(\mathbb{E}[\partial_{x}f(\bm{\xi},x)],\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x)\Big)=\sup_{|w|=1}\Big|\mathbb{E}[f_{x}^{\prime}(\bm{\xi},x;w)]-\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f_{x}^{\prime}(\bm{\xi}^{i},x;w)\Big|.$
Let $P=\mathbb{P}\circ\bm{\xi}^{-1}$. It suffices to show that, for any fixed $w\in\{-1,1\}$, the class of Borel measurable functions $\mathcal{S}=\{f_{x}^{\prime}(\cdot,x;w)\mid x\in\mathbb{R}\}$ is $P$-Glivenko–Cantelli; i.e.,
@@
which is well defined as explained below. For simplicity, assume $w=1$. Using the fact that $f(\xi,\cdot)$ is $L(\xi)$-Lipschitz, we have $|f_{x}^{\prime}(\xi,x;w)|\leq L(\xi)$ for any $x$. Because $\mathbb{E}[L(\bm{\xi})]<\infty$, the class $\mathcal{S}$ admits an integrable envelope $L$. From [25, Theorem 7.42] and [18, Theorem 24.1], the function $t\mapsto f_{x}^{\prime}(\xi,t;w)$ is non-decreasing and right-continuous (when w=1w=-1, it is non-increasing and left-continuous), thereby $\mathcal{S}$ is pointwise measurable [28, Example 2.3.4] and hence $P$-measurable [28, Definition 2.3.3] with (7) being well defined. With monotonicity and boundedness of $f_{x}^{\prime}(\xi,\cdot;w)$, set
@@
where $l,u:\Xi\to\mathbb{R}$ are Borel measurable with $\mathbb{E}[u(\bm{\xi})]-\mathbb{E}[l(\bm{\xi})]\leq 2\mathbb{E}[L(\bm{\xi})]$. If $\mathbb{E}[u(\bm{\xi})]=\mathbb{E}[l(\bm{\xi})]$, then $f_{x}^{\prime}(\xi,\cdot;w)$ is almost surely constant and the claim is trivial. Assume $\mathbb{E}[(u-l)(\bm{\xi})]>0$, and fix any $0<\varepsilon<\tfrac{2}{3}\mathbb{E}[(u-l)(\bm{\xi})]$. For each $n\in\mathbb{N}$, define
@@
where the attainment is from the right-continuity of $t\mapsto\mathbb{E}[f_{x}^{\prime}(\bm{\xi},t;w)]$. Let $N\in\mathbb{N}$ be the first number such that $\mathbb{E}[f_{x}^{\prime}(\bm{\xi},t_{N};w)]\geq\mathbb{E}[u(\bm{\xi})]-\tfrac{\varepsilon}{2}$. By construction, we have $N\leq 1+2\mathbb{E}[L(\bm{\xi})]/\varepsilon<\infty$ and $\mathbb{E}[|u(\bm{\xi})-f_{x}^{\prime}(\bm{\xi},t_{N};w)|]\leq\varepsilon$. By [18, Theorem 24.1], we get $f_{x}^{\prime}(\xi,t_{n};-w)=\lim_{s\uparrow t_{n}}f_{x}^{\prime}(\xi,s;w)$. Using dominated convergence theorem and the definition of $t_{n}$, for any $n\in\mathbb{N}$, one has
@@
which yields that $\mathbb{E}[|-f_{x}^{\prime}(\bm{\xi},t_{n+1};-w)-f_{x}^{\prime}(\bm{\xi},t_{n};w)|]\leq\varepsilon$ whenever $t_{n+1}\neq t_{n}$ and $\mathbb{E}[|-f_{x}^{\prime}(\bm{\xi},t_{1};-w)-l(\bm{\xi})|]\leq\varepsilon$. Then, the following (possibly empty when $t_{n}=t_{n+1}$) sets of functions
@@
Hence, the bracketing number $\mathcal{N}_{[\,]}(\varepsilon,\mathcal{S},L^{1}(P))$ is finite. Invoking [28, Theorem 2.4.1], we conclude that the class $\mathcal{S}$ is $P$-Glivenko–Cantelli. Applying the same argument to w=1w=-1, with simple adjustments for nonincreasingness and left-continuity, and then taking the maximum over $w\in\{-1,1\}$ completes the proof.
@@ 3.4 Proof of Theorem 3
$f(\xi,x)=h(g(\xi,x))+35\|x\|^{2}=\max\{g(\xi,x),0\}+35\|x\|^{2};$
see Figure 2. Define $\delta^{\nu}=(2240\lceil 2^{\nu+1}\ln(\nu+1)\rceil^{4})^{-1}.$
@@
$|g(\xi,x)-g(\xi,y)|\leq|x_{1}-y_{1}|+\mathop{\sum}\nolimits_{j\in\{k_{1},k_{2}\}}|\psi_{j}(x_{2})-\psi_{j}(y_{2})|\leq 61\|x-y\|.$
Then $g(\xi,\cdot)$ is $C$-Lipschitz continuous for every $\xi\in\Xi$. We next show that $g(\xi,(0,\cdot))$ is differentiable at 0. If $t\in\operatorname{supp}(\psi_{k})$ for some $k$, then $t\geq\frac{1}{2k}$ and $|g(\xi,(0,t))|\leq\eta_{k}\leq t^{2}.$ If $t\notin\cup_{k=1}^{\infty}\operatorname{supp}(\psi_{k})$, then $g(\xi,(0,t))=0$. Hence $|t|^{-1}|g(\xi,(0,t))-g(\xi,(0,0))|\leq|t|\to 0$ as $t\to 0$, so $g(\xi,(0,\cdot))$ is differentiable at 0 with $g(\xi,(0,\cdot))^{\prime}(0)=0$. Consequently, a similar argument based on disjoint supports gives
$\nabla_{x}g(\xi,x)=\left[\begin{array}[]{c}1\\ \mathop{\sum}\nolimits_{k=1}^{\infty}\psi_{k}^{\prime}(x_{2})(2\mathop{\textnormal{bit}}\nolimits_{k}(\xi)-1)\end{array}\right]$
and $\|\nabla_{x}g(\xi,x)-\nabla_{x}g(\xi,y)\|\leq C\|x-y\|.$ Hence $g(\xi,\cdot)$ is $C$-smooth. Hence, $f(\xi,\cdot)$ is convex and $2C$-Lipschitz on $X$. From Lemma 1, the functions $\xi\mapsto\mathop{\textnormal{bit}}\nolimits_{k}(\xi)$, $\xi\mapsto g(\xi,x)$, and then $\xi\mapsto f(\xi,x)$ are measurable. Therefore, $f$ and $g$ are Carathéodory. Moreover, for any $x\in\mathbb{R}^{2}$, one has $|\mathbb{E}[f(\bm{\xi},x)]|\leq\mathbb{E}[|g(\bm{\xi},x)|]+35\|x\|^{2}\leq|x_{1}|+\tfrac{1}{64}+35\|x\|^{2}<\infty.$
@@
and $f(\xi,\cdot)$ is smooth on $\cup_{k}\mathbb{B}(p_{k},\Delta_{k})$ for any $\xi\in\Xi$. Hence, for any $k\in\mathbb{N}$ and $y\in\mathbb{B}(p_{k},\Delta_{k})$, by [19, Exercise 10.26], we can write
@@
Let $\mathcal{F}$-measurable $E\subset\Omega$ with $\mathbb{P}(E)=1$, $\bar{\bm{\nu}}$, and $\{\bm{k}^{\nu}\}_{\nu}$ be defined in Lemma 2. For any $\omega\in E$, $\nu\geq\bar{\bm{\nu}}(\omega)$, one has $\omega\in E^{\nu}_{\bm{k}^{\nu}(\omega)}$ for some $\nu\leq\bm{k}^{\nu}(\omega)\leq\lceil 2^{\nu+1}\ln(\nu+1)\rceil$. Then, $\mathop{\textnormal{bit}}\nolimits_{\bm{k}^{\nu}(\omega)}(\bm{\xi}^{i}(\omega))=1$ for all $i\in[\nu]$. Let $\bm{p}^{\nu}(\omega)=p_{\bm{k}^{\nu}(\omega)}\in\{p_{k}\}_{k}$. For any $y\in\mathbb{B}(\bm{p}^{\nu}(\omega),\Delta_{\bm{k}^{\nu}(\omega)})$, we have
@@
$0\leq\delta^{\nu}=\big(2240\lceil 2^{\nu+1}\ln(\nu+1)\rceil^{4}\big)^{-1}=\Delta_{\lceil 2^{\nu+1}\ln(\nu+1)\rceil}\leq\Delta_{\bm{k}^{\nu}(\omega)}.$
Therefore, for any $\omega\in E$, we conclude that $\bm{p}^{\nu}(\omega)\in\{p_{k}\mid\nu\leq k\leq\lceil 2^{\nu+1}\ln(\nu+1)\rceil\}\subset D^{\nu}$ and
@@
3.5 Proof of Corollaries 3 and 4
We use the notation in Section 3.2. The proof then follows from the following three observations.
  1. (a)
    Let $h(\xi,x)=f(\xi,x)-x$. For any $\omega\in E$ and large $\nu\in\mathbb{N}$, the point $\bm{p}^{\nu}(\omega)$ in Step 4 of Section 3.2 is a local minimizer of $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}h(\bm{\xi}^{i}(\omega),\cdot)$ and $\bm{p}^{\nu}(\omega)\to 0$ as $\nu\to\infty$.
  2. (b)
    For $\omega\in E$ and any $\nu\in\mathbb{N}$, one has $1\in\partial\left(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\bm{\xi}^{i}(\omega),\cdot)\right)(0)$.
  3. (c)
    $\partial\mathbb{E}[f(\bm{\xi},\cdot)](0)\subset[-\frac{1}{2},\frac{1}{2}]$, hence $0\notin\partial\mathbb{E}[h(\bm{\xi},\cdot)](0)\subset[-\frac{3}{2},-\frac{1}{2}]$.
We now prove these claims. From Steps 3 and 5 in Section 3.2, for any $\nu\geq\bar{\bm{\nu}}(\omega)$, we know that $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}h(\bm{\xi}^{i}(\omega),\cdot)$ is constant on $\mathbb{B}(\bm{p}^{\nu}(\omega),\delta^{\nu})$ with $\delta^{\nu}>0$. Moreover, $|\bm{p}^{\nu}(\omega)|\leq p_{\nu}=1/\nu$, hence $\bm{p}^{\nu}(\omega)\to 0$. This proves (a). Next, for any fixed $\nu$, we consider the sequence $\{\bm{p}^{\tau}(\omega)\}_{\tau\geq\max\{\nu,\bar{\bm{\nu}}(\omega)\}}$. By Step 4 in Section 3.2 and $\nu\leq\tau$, one has $\mathop{\textnormal{bit}}\nolimits_{\bm{k}^{\tau}(\omega)}(\bm{\xi}^{i}(\omega))=1$ for all $i\in[\nu]$. Hence, we have $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\nabla_{x}f(\bm{\xi}^{i}(\omega),\bm{p}^{\tau}(\omega))=1$. Together with $\bm{p}^{\tau}(\omega)\to 0$ as $\tau\to\infty$, this implies $1\in\partial(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\bm{\xi}^{i}(\omega),\cdot))(0)$, which establishes (b). Finally, using Step 2 in Section 3.2 and Fubini, we obtain $\mathbb{E}[f(\bm{\xi},x)]=\tfrac{1}{2}\int_{0}^{\max\{x,0\}}\mathds{1}_{\cup_{k}B_{k}}(t)\textrm{d}t.$ Therefore, the map $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is $\tfrac{1}{2}$-Lipschitz continuous, and thus $\partial\mathbb{E}[f(\bm{\xi},\cdot)](0)\subset[-\frac{1}{2},\frac{1}{2}]$, proving (c). In summary, Corollary 3 follows from (a) and (c), while Corollary 4 follows from (b) and (c).
3.6 Details of the DC Construction
Let $\bm{\xi}=(\bm{\xi}_{1},\bm{\xi}_{2})$ be defined by iid random variables $\bm{\xi}_{1},\bm{\xi}_{2}$ on $(\Omega,\mathcal{F},\mathbb{P})$ with uniform distributions on $\Xi=[0,1]$. We begin with a new lemma, whose proof is a straightforward modification of Lemma 2.
Lemma 3.
There exist an $\mathcal{F}$-measurable set $E\subset\Omega$ with $\mathbb{P}(E)=1$ and $\mathbb{N}$-valued random variables $\bar{\bm{\nu}},\bm{k}^{1},\bm{k}^{2},\ldots$ such that for any $\omega\in E$ and any $\nu\geq\bar{\bm{\nu}}(\omega)$, one has $\nu\leq{\bm{k}^{\nu}(\omega)}\leq\lceil 4^{\nu+1}\ln(\nu+1)\rceil$,
$\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}^{1}_{1}(\omega))=\cdots=\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}_{1}^{\nu}(\omega))=1,\quad\text{and}\quad\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}^{1}_{2}(\omega))=\cdots=\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}_{2}^{\nu}(\omega))=0.$
Consider the setting and analysis in Section 3.4, replacing Lemma 2 with Lemma 3 wherever appropriate. Using $g:\Xi\times\mathbb{R}^{2}\to\mathbb{R}$ in Section 3.4, define a DC function $\ell:\Xi^{2}\times\mathbb{R}^{2}\to\mathbb{R}$ as
$\ell(\xi,x)=\max\{g(\xi_{1},x),0\}+35\|x\|^{2}-\max\{g(\xi_{2},x),0\}-35\|x\|^{2}.$
Let $h(\xi,x)=\ell(\xi,x)-x_{1}$. Parallel to Section 3.5, we observe:
  1. (a)
    For any $\omega\in E$ and large $\nu\in\mathbb{N}$, the point $\bm{p}^{\nu}(\omega)$ in Step 4 of Section 3.4 using Lemma 3 instead of Lemma 2 is a local minimizer of $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}h(\bm{\xi}^{i}(\omega),\cdot)$ and $\bm{p}^{\nu}(\omega)\to 0$ as $\nu\to\infty$.
  2. (b)
    For $\omega\in E$ and any $\nu\in\mathbb{N}$, one has $(1,0)\in\partial\left(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\ell(\bm{\xi}^{i}(\omega),\cdot)\right)(0)$.
  3. (c)
    $\partial\mathbb{E}[\ell(\bm{\xi},\cdot)](0)=\{0\}$, hence $0\notin\partial\mathbb{E}[h(\bm{\xi},\cdot)](0)=\{(-1,0)\}$.
We now prove these claims. From Step 3 in Section 3.4, for any $\nu\geq\bar{\bm{\nu}}(\omega)$ and $i\in[\nu]$, one has $g(\bm{\xi}^{i}_{1}(\omega),\bm{p}^{\nu}(\omega))=\bm{p}_{1}^{\nu}(\omega)+\eta_{\bm{k}^{\nu}(\omega)}>0$ and $g(\bm{\xi}^{i}_{2}(\omega),\bm{p}^{\nu}(\omega))=\bm{p}_{1}^{\nu}(\omega)-\eta_{\bm{k}^{\nu}(\omega)}<0.$ Therefore, for any $y$ near $\bm{p}^{\nu}(\omega)$, one has $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}h(\bm{\xi}^{i}(\omega),y)=\eta_{\bm{k}^{\nu}(\omega)},$ which is a constant. Moreover, $\|\bm{p}^{\nu}(\omega)\|\leq 1/\nu$, hence $\bm{p}^{\nu}(\omega)\to 0$. This proves (a). Next, for any fixed $\nu$, we consider the sequence $\{\bm{p}^{\tau}(\omega)\}_{\tau\geq\max\{\nu,\bar{\bm{\nu}}(\omega)\}}$. By Step 4 in Section 3.4 (using Lemma 3) and $\nu\leq\tau$, for all $i\in[\nu]$, one has $\mathop{\textnormal{bit}}\nolimits_{\bm{k}^{\tau}(\omega)}(\bm{\xi}_{1}^{i}(\omega))=1$ and $\mathop{\textnormal{bit}}\nolimits_{\bm{k}^{\tau}(\omega)}(\bm{\xi}_{2}^{i}(\omega))=0.$ Hence, from Step 3 in Section 3.4, we have $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\nabla_{x}\ell(\bm{\xi}^{i}(\omega),\bm{p}^{\tau}(\omega))=(1,0)$. Together with $\bm{p}^{\tau}(\omega)\to 0$ as $\tau\to\infty$, this establishes (b). Finally, since $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$ have identical distributions, observe that $\mathbb{E}[\ell(\bm{\xi},x)]\equiv 0$. Therefore, $\partial\mathbb{E}[\ell(\bm{\xi},\cdot)](0)=\{0\}$, proving (c).